OpenStudy (anonymous):

given the function f(x)= 0.5(3)^x, what is the value of f^-1(7)

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@campbell_st I opened a new one

OpenStudy (anonymous):

would I have to do 0.5 times itself 3 times?

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@kjones0331

OpenStudy (anonymous):

@RadEn @SithsAndGiggles

OpenStudy (anonymous):

\[f(x)=y=\frac{1}{2}(3)^x\] Solve for \(x\) in terms of \(y\): \[2y=3^x\\ \ln(2y)=\ln 3^x\\ \ln(2y)=x\ln 3\\ x=\frac{\ln(2y)}{\ln3}\\ x=\log_3(2y)~~\Rightarrow~~f^{-1}(x)=\log_3(2x)\] Now plug in \(x=7\). Or, more simply, just solve the following equation for \(x\): \[7=\frac{1}{2}(3)^x\]

OpenStudy (anonymous):

hmm ok so I subtract 7 from both sides

OpenStudy (anonymous):

or divide 7 I believe

OpenStudy (anonymous):

is it division @SithsAndGiggles

OpenStudy (anonymous):

No, that's not how you would solve for \(x\). One of the first steps would be to take it down from the exponent: \[7=\frac{1}{2}(3)^x\\ 2\cdot7=2\cdot\frac{1}{2}(3)^x\\ 14=3^x\] To lower the exponent, you take the logarithm of both sides: \[\ln14=\ln3^x\] One of the properties of logarithms says \(\ln a^b=b\ln a\), which gives you \[\ln14=x\ln3\] Then, \[\frac{\ln14}{\ln3}=\frac{x\ln3}{\ln3}\\ \frac{\ln14}{\ln3}=x\] Another property of logarithms: \(\dfrac{\ln a}{\ln b}=\log_ba\), so, \[x=\log_3(14)\approx\cdots\]

OpenStudy (anonymous):

I got 4.67

OpenStudy (anonymous):

No, that's not right...

OpenStudy (anonymous):

2.402?

OpenStudy (anonymous):

@SithsAndGiggles heloo

OpenStudy (anonymous):

Yes, that's the right answer. Sorry